A bullet after penetrating 3cm of a wall loses half of its speed. How far will the bullet penetrate the wall afterwards?


1
Expert's answer
2021-01-30T06:26:26-0500

Loss of kinetic energy:

"\frac{mv^2}{2}-\frac{m(\frac v2)^2}{2}=Fx," where "x=3\cdot 10^{-2},"

"F=\frac{\frac{mv^2}{2}}{4\cdot10^{-2}}."

Total:

"\frac{mv^2}{2}=F\Delta x,"

"\Delta x=\frac{\frac{mv^2}{2}}{F}=4\cdot 10^{-2},"

"\Delta x=x+x_1,"

"x_1=\Delta x-x=1\cdot 10^{-2}=1~cm."